Help on my Finance question please.

Integrative: Risk and valuation Hamlin Steel Company wishes to determine the value of Craft Foundry, a firm that it is considering acquiring for cash. Hamlin wishes to determine the applicable discount rate to use as an input to the constant-growth valuation model. Craft’s stock is not publicly traded. After studying the required returns of firms similar to Craft that are publicly traded, Hamlin believes that an appropriate risk premium on Craft stock is about 9%. The risk-free rate is currently 5%. Craft’s dividend per share for each of the past 6 years is shown in the following table.

Year Dividend per share
2019 $3.44  
2018 3.28
2017 3.15
2016 2.90
2015 2.75
2014 2.45
  1. Given that Craft is expected to pay a dividend of $3.68 next year, determine the maximum cash price that Hamlin should pay for each share of Craft.

  2. Describe the effect on the resulting value of Craft of

    1. A decrease in its dividend growth rate of 2% from that exhibited over the 2014–2019 period.

    2. A decrease in its risk premium to 4%.
Answer #1

(a): Required return on Craft’s stock = 9%+5% = 14%

The maximum cash price that Hamlin should pay for each share of Craft is:

First we will determine the CAGR growth rate of the dividends. Thus CAGR = (3.44/2.45)^(1/5) – 1 = 7.02%

Thus price = dividend/r-g

= 3.68/(14%-7.02%)

= $52.75

(b-1): Here g will become 7.02-2 = 5.02%. Also next dividend will be = 3.44*(1+2%) = 3.5088

Thus price = 3.5088/(14%-5.02%)

= $39.09

Thus per share value will decline to $39.09 in this case.

(b-2): Here required return will be = 4%+5% = 9%

Thus price = 3.68/9%-7.02%

= $186.17

Thus per share value will increase to $186.17 in this case.